​(Calculating project cash flows and​ NPV)  The Chung Chemical Corporation is considering the purchase of a chemical analysis machine. Although the machine being considered will result in an increase in earnings before interest and taxes of $ 33000 per​ year, it has a purchase price of ​$115 000​, and it would cost an additional ​$6 000 after tax to correctly install this machine. In​ addition, to properly operate this​ machine, inventory must be increased by ​$5 500. This machine has an expected life of 10 ​years, after which it will have no salvage value. ​ Also, assume simplified​ straight-line depreciation, that this machine is being depreciated down to​ zero, a 32 percent marginal tax​ rate, and a required rate of return of 7 percent.

a.  What is the initial outlay associated with this​ project?

b.  What are the annual​ after-tax cash flows associated with this project for years 1 through 9​?

c.  What is the terminal cash flow in year 10 ​(that is, the annual​ after-tax cash flow in year 10 plus any additional cash flows associated with the termination of the​ project)?

d.  Should this machine be​ purchased?
Answer #1

a) Initial Outlay = Investment + NWC = 115,000 + 6,000 + 5,500 = 126,500

b) After-tax cash flow = EBIT x (1 - tax) + Depreciation = 33,000 x (1 - 32%) + 121,000 / 10 = 34,540

c) In year 10, working capital will be recovered, cash flow = 34,540 + 5,500 = 40,040

d) Rate of return of the project can be calculated using I/Y function on a calculator

N = 10, PV = -126,500, PMT = 34,540, FV = 5,500 => Compute I/Y = 24.33% is the IRR of the project

As IRR > 7%, the machine should be purchased.